### v1

# CITADEL: A Semi-Supervised Active Learning Framework for Malware Detection Under Continuous Distribution Drift

Abstract
Android malware evolves rapidly, leading to concept drift that degrades the performance of traditional machine learning (ML)-based detection systems. While recent approaches incorporate active learning and hierarchical contrastive loss to handle this drift, they remain fully supervised, computationally expensive, and perform poorly on real-world datasets with long temporal spans. In particular, our evaluation highlights these limitations, particularly on LAMDA, a 12-year longitudinal dataset exhibiting substantial distributional shifts. Moreover, manual expert labeling cannot scale with the daily emergence of over 450,000 new malware samples, leaving most samples unlabeled and underutilized.
To address these challenges, we propose CITADEL, a robust semi-supervised active learning framework for Android malware detection. To bridge the gap between image-domain semi-supervised learning and binary feature representations of malware, we introduce malware-specific augmentations, Bernoulli bit flips and masking, that simulate realistic drift behaviors. CITADEL further integrates supervised contrastive loss to improve boundary sample discrimination and combines it with a multi-criteria active learning strategy based on prediction confidence, -norm distance, and boundary uncertainty, enabling effective adaptation under limited labeling budgets. Extensive evaluation on four large-scale Android malware benchmarks—APIGraph, Chen-AZ, MaMaDroid, and LAMDA demonstrates that CITADEL outperforms prior work, achieving F1 score of over 1%, 3%, 7%, and 14% respectively, using only 40% labeled samples. Furthermore, CITADEL shows significant efficiency over prior work incurring faster training and fewer operations.
Availability. The code is available at https://github.com/IQSeC-Lab/CITADEL.git.
I Introduction
The Android ecosystem, with over 3 billion active devices, has become a primary target for malware developers [1, 2]. The AV‑TEST Institute reports around 5 million new Android malware and potentially unwanted applications (PUAs) annually [3], posing a growing challenge to reliable threat detection. Traditional methods such as sandbox analysis and signature matching are labor-intensive, brittle to obfuscation, and struggle to scale [4, 5]. As a scalable alternative, machine learning (ML)-based detection systems have gained widespread adoption in both academia and industry [6, 7, 8], including integration by vendors such as Avast [9]. These systems learn from structural and behavioral features, and have demonstrated strong performance and generalization [10, 11].
A core challenge in ML-based malware detection is concept drift–the change in the statistical properties of malware and benign samples over time. State-of-the-art concept drift adaptation method leverage active learning (AL) with hierarchical contrastive learning and uncertainty-based sample selection [11]. Though effective on short-span benchmarks—yielding 92.4% F1 on the 7-year APIGraph dataset [13], its performance drops sharply on longer temporal spans, dropping to 43% on the 12-year LAMDA benchmark [15] and further to 30% on a 12-year curated dataset derived from MaMaDroid [14] feature (see Figure 1).
Furthermore, with over 450,000 new malware samples emerging daily [16], producing high-quality labels for these enormous volume is time, resource, and budget intensive; human analysts can review only around 80 samples per day [17], and vendors may take weeks to agree on newly discovered variants [18, 19]. Existing concept drift adaptation methods [20, 11] rely heavily on supervised learning, assuming continuous access to labeled data and focusing primarily on short-term evaluations. Such assumptions do not hold in realistic, long-term drift scenarios, where labeled data are scarce and the malware distribution evolves unpredictably. These challenges highlight the need for semi-supervised learning (SSL), which can leverage the abundance of unlabeled samples to maintain adaptability and performance in dynamic, label-scarce environments [21].
To address long-term malware concept drift adaptation under limited labeling budgets, we propose CITADEL, a SSL framework adapted from FixMatch [22], specifically tailored for Android malware detection under distribution shift. Although prior work utilize FixMatch [12], their goal is to tackle label noise and data imbalance, which does not address the challenges of concept drift. Moreover, FixMatch is originally designed for computer vision domain and it suffers significantly to offer even reasonable performance for malware detection (see Table III). As such, we improve it for malware feature representations with domain-relevant transformations—Bernoulli bit flip and Bernoulli feature mask.
In practice, there is no deterministic pattern describing how feature distributions evolve over time [7], as new malware families may stop using certain API calls, permissions, or intent filters, while adopting others. As such, we hypothesize that Bernoulli-based transformations would be better fit to model the unpredictable nature of adversary induced distribution shifts in malware evolution. To simulate the uncertainty of adversary induced drift, we adopt a probabilistic approach using the Bernoulli distribution to randomly select features for modification. In Bernoulli bit flip, a randomly chosen subset of binary features (such as API calls or permissions) is flipped turning a previously used feature “off” or an unused one “on.” This simulates behavioral changes where malware starts or stops using specific functionalities. For Bernoulli feature mask deactivates a randomly selected subset of features, representing the disappearance of certain patterns (such as deprecated API calls or obsolete permissions) in newer malware variants. Since static features cannot capture entirely new APIs introduced in the future, this feature-space perturbation provides a principled way to simulate evolving distributions.
Our analysis show that CITADEL learns feature representations that effectively capture temporal drift. Adapting to such evolving patterns requires AL to selectively query the most informative samples [11]. Furthermore, malware drift is often abrupt and unpredictable, making such ordered progression ineffective. Many drifted samples tend to lie near the decision boundary—regions where the model is most uncertain and prone to misclassification [23, 24]. Additionally, outlier samples often appear as distant points in the latent space, further challenging accurate classification [20]. To address these issues, we propose a multi-criteria sample selection strategy that prioritizes: (i) low-margin samples (small confidence differences between top predictions), (ii) samples with high -norm distances, and (iii) low-confidence predictions. Together, these criteria target samples that are both ambiguous and influential for improving decision boundaries.
Moreover, our analysis shows that boundary samples where benign and malware features overlap are particularly challenging to classify. To better handle such drifted and ambiguous cases, we introduce an auxiliary objective that promotes better separation of such ambiguous cases, complementing the semi-supervised loss. Figure 3 illustrates the overall architecture of CITADEL. Our extensive evaluation shows that CITADEL attains strong performance, achieving F1 scores of 93.5% on APIGraph, 82.7% on Chen-AZ, 44.9% on MaMaDroid, and 77.7% on LAMDA. In addition to accuracy, CITADEL provides substantial efficiency gains, offering up to 24 faster training and requiring 13 fewer operations than prior methods.
In summary, the contributions of this work are as follows:
We propose CITADEL, a SSL framework for Android malware detection under long-term concept drift. CITADEL introduces malware domain-specific augmentations, Bernoulli bit flip and Bernoulli feature mask.
To better discriminate ambiguous boundary samples, we integrate a supervised contrastive loss [25] alongside the standard SSL objective, promoting clearer separation of low-margin samples from different classes in latent space.
We propose a Multi-Criteria sample selection strategy for AL that selects highly informative samples by jointly considering boundary margin, -norm distance, and low-confidence predictions—enabling efficient adaptation under labeling budget constraints.
We evaluate CITADEL on four Android malware benchmarks—APIGraph [13], Chen-AZ [11], MaMaDroid [14], and LAMDA [15]—under static and AL settings. CITADEL consistently outperforms prior methods, achieving F1 gains of on APIGraph, on Chen-AZ, on MaMaDroid, and on LAMDA. It also offers up to faster training and requires fewer operations, making it effective and efficient for deployment.
II Background and Related Work
In this section, we outline prior research on handling concept drift in malware analysis, including detection, adaptation, and learning strategies such as semi-supervised, active, curriculum, and continual learning.
Concept Drift Detection
Detecting concept drift is crucial for maintaining the long-term effectiveness of malware classifiers as threats evolve. Strategies in incremental learning [26, 27] aim to maintain accuracy while minimizing retraining. Key strategies include Gama et al. [28]’s drift detection methods to track model performance, Yang et al. [20]’s CADE framework using contrastive learning for dimensionality reduction and drift detection, and Barbero et al. [29]’s TRANSCENDENT framework that uses statistical analysis to quantify prediction credibility. OWAD [30] proposes an unsupervised method for detecting normality shift—a form of concept drift in zero-positive anomaly detection where only normal data is available. These methodologies enable reliable adaptation and decision-making without frequent retraining.
Concept Drift Adaptation
Upon detecting concept drift, effective adaptation is crucial for maintaining classifier performance against evolving malware and goodware (benign). Several prior works have jointly addressed both drift detection and adaptation. Jordaney et al. [31] developed a framework for early adaptation using kernel-based drift detection. Mishra et al. [32] used clustering to find feature shifts in malware. Chen-AL [11] applied hierarchical contrastive and active learning for efficient adaptation, lowering labeled data needs while improving F1 scores. Domain adaptation methods has also been studied for drift adaptation. Specifically, Pan et al. [26] started this, followed by Li et al. [27], who used graph-based techniques to focus on drift-invariant features resilient to obfuscation. These efforts underscore the importance of drift detection and adaptation for reliable malware classification. On the other hand, OWAD [30] selects key drift-inducing samples via an optimization-based explainer, reducing labeling cost. It updates the model using parameter-wise regularization that preserves important knowledge while learning new patterns.
Semi-Supervised Learning (SSL)
SSL has become increasingly valuable for malware detection, where labeled samples are limited but unlabeled data are abundant. Prior studies have shown that SSL can enhance detection accuracy while reducing manual effort. DeepReflect [33] integrates analyst feedback to minimize reverse engineering, while Kan et al. [34] demonstrate that pseudo-labeling helps classifiers adapt to new threats without frequent retraining. Apruzzese et al. [21] show that leveraging unlabeled data improves detection across diverse malware families. Collectively, these works emphasize SSL’s potential to build scalable and adaptive malware detection systems.
SSL has also gained traction in computer vision (CV) due to abundant unlabeled data and the high cost of manual annotation. Frameworks such as FixMatch [22], FlexMatch [35], FreeMatch [36], SoftMatch [37], and Dash [38] follow a common pipeline—generating pseudo-labels from weakly augmented inputs, filtering them by confidence, and enforcing consistency on strongly augmented variants. Each method refines this process through different confidence strategies, including class-wise, sample-wise, or momentum-based adjustments. However, image augmentations like flips or color shifts are inapplicable to binary malware features.
Active Learning in Malware Detection
Active learning (AL) is used in security to reduce labeling costs and keep models adaptable. Instead of labeling all data, AL queries uncertain samples for annotation, aiding rapid retraining with new malware. This is useful in malware detection, where expert labeling is costly and malware evolves quickly. Fang et al.’s BODMAS framework [39] found uncertainty sampling boosts malware detection performance. Zhou et al.’s APIGraph [40] enhanced this with meta-features to merge similar API calls, improving Android malware detection’s robustness with AL.
Other learning paradigms have been explored for malware classification. Casandra [41] uses online learning, updating the classifier with each newly labeled app for continual adaptation—though this is computationally demanding and impractical in resource-limited settings. DroidEvolver [42] applies self-training, assigning pseudo-labels to unlabeled data to reduce manual effort; however, Demetrio et al. [43] warn this can cause self-poisoning, where model errors are reinforced.
Curriculum Learning (CurL)
CurL primarily focuses on structuring training from easy to hard examples to improve model stability and convergence. Zhou et al. [44] adjusts sample difficulty based on model uncertainty, allowing the curriculum to evolve during training. Shashidhar et al. [45] present a model curriculum for teaching malware analysis, combining theoretical foundations with practical exercises such as reverse engineering and dynamic analysis which is out of scope in this paper.
Continual Learning (CL)
Unlike active learning, which aims to reduce annotation costs by selecting informative samples, CL focuses on maintaining performance across evolving tasks or data distributions by mitigating catastrophic forgetting [46]. Rahman et al.[47] first demonstrated that CL methods developed for computer vision (CV) perform poorly in the malware domain, highlighting the need for a domain specific CL framework. Recently, MalCL[48] proposes a GAN-based Class-IL framework that leverages feature matching loss and intermediate-layer-guided replay sample selection to improve generative replay and reduce forgetting. Furthermore, MADAR [49] proposes a diversity-aware replay strategy that balances representative and anomalous samples, capturing both inter- and intra-family variations to improve performance.
III Preliminaries
III-A Problem Setup and Notations
We formulate our task as a semi-supervised malware detection problem under concept drift, where the data distribution of the malware (or benign) samples evolves over time. Let denote the input space of feature representations extracted from Android application (e.g., API call sequences or DREBIN features [6]), where 0 corresponds to absence and 1 corresponds to the presence of the feature of a sample; and let denote the label space, where corresponds to benign software and to malware.
Concept drift, arising from the evolving behavior of malware over time. We illustrate this in Figure 2 using the t-SNE projections latent space representations of various training samples of the APIGraph datasets that are spanned across several months. Specifically, some malware samples align closer to benign samples, and vice versa, leading to misclassifications of both classes. Additionally, certain samples deviate significantly from both the classes, forming outliers. Furthermore, boundary samples are particularly susceptible to shifts in data distribution and may challenge classifier robustness.
At each time step , the detector receives a small labeled dataset and a large pool of unlabeled data , where . The goal is to learn a malware detector with parameters that estimates the posterior probability under the current data distribution , such that the detection performance generalizes well even as shifts over time due to concept drift. Concept drift refers to a temporal shift in the joint distribution such that:
This change can be caused by new malware variants, evasion techniques, or benign software mimicking malicious behavior and vice versa. Drift can affect the input distribution (covariate drift), the conditional distribution (real concept drift), or both. To address evolving distributions, the learning framework must support: (1) adaptive updates to the model without full retraining, (2) robust use of unlabeled data under shifting domains, and (3) efficient selection of informative samples when labeling budgets are constrained.
III-B Motivating Example
A core challenge in malware detection is the emergence of new variants that introduce previously unseen or rare semantic features, leading to concept drift. Traditional supervised learning models are typically trained on static, labeled datasets and often fail to generalize when confronted with novel traits at test time. Compounding this issue, producing high-quality labels for malware is both time-consuming and resource-intensive—prior work shows that vendors can take weeks to reach consensus on newly discovered variants [18, 19]. Moreover, malware analysts can manually review only about 80 samples per day [17], leaving the vast majority of malware unlabeled. This creates an untapped opportunity to utilize the abundance of unlabeled samples in the detection pipeline [21]. Motivated by this observation, we explore SSL as a means to both detect and adapt to concept drift by jointly leveraging labeled and unlabeled data. In particular, we investigate the FixMatch framework [22], which has demonstrated strong performance in other domains by effectively utilizing unlabeled samples.
FixMatch combines pseudo-labeling with consistency regularization through weak and strong data augmentation techniques. For an unlabeled input sample , FixMatch generates a pseudo-label by applying a weak augmentation and passing it through the current model. This pseudo-label is then used as a supervisory signal for the corresponding strongly augmented version . The learning objective enforces consistency between the model’s predictions for and , thereby encouraging the model to produce stable predictions under input perturbations. Weak augmentations (e.g., horizontal flips, minor translations, rotations) preserve the semantic content of the input and facilitate reliable pseudo-labeling. Strong augmentations (e.g., color jitter, brightness alteration, cutout) introduce stochastic feature perturbations that can mask or distort key input features [50]. Despite this, the model is trained to maintain prediction consistency, thereby forcing it to generalize beyond specific feature dependencies.
In the context of malware detection, this mechanism is particularly useful for learning from drift-induced features. Let be a malware sample containing both a known semantic feature and a novel, previously unseen feature . The weak augmentation retains both features, enabling the model to generate a high-confidence pseudo-label based on . If the strong augmentation masks via stochastic feature perturbation, the detector must rely on to maintain consistency with , thereby incorporating into the learned decision boundary. This implicit adaptation supports concept drift handling by progressively integrating novel features without explicit annotation. Prior work demonstrates that this mechanism promotes the learning of low-density decision boundaries and improved feature alignment between labeled and unlabeled samples [50].
In our setting, FixMatch enables the model to bridge concept drift by exploiting the structure of unlabeled data, thereby facilitating more robust detection of unseen or evolving malware variants. The training dataset at step in the Figure 3 is divided into two sets labeled, at step and unlabeled at step . Let denote the labeled dataset and the unlabeled dataset. The overall FixMatch objective is a weighted sum of the supervised loss on labeled data and the unsupervised consistency loss on high-confidence pseudo-labeled data:
where is the standard cross-entropy (CE) loss computed over labeled samples:
and is the unsupervised loss enforcing consistency on strongly augmented unlabeled inputs, conditioned on high-confidence pseudo-labels:
Here, denotes the malware detector parameterized by , is the strongly augmented view of the unlabeled input , and is the predicted softmax probability from the weakly augmented version used to obtain pseudo-label . The indicator function filters pseudo-labels based on a confidence threshold . The hyperparameter controls the influence of the unsupervised loss. In our experiments, we set and , following best practices from prior work [22]. This formulation encourages the model to produce consistent predictions across different input perturbations and to generalize from labeled data while gradually adapting to new or shifting semantic features, capabilities essential for effective malware detection under concept drift.
IV Overview of CITADEL
In this section, we introduce our semi-supervised active learning (AL) framework, CITADEL, for adapting to concept drift in malware detection. The overall process is outlined in Algorithm 1 (Semi-supervised Initial Training) and Algorithm 2 (Active Learning).
We first propose two domain-specific augmentations—Bernoulli bit flip and feature mask—to emulate realistic malware evolution and perturb semantic features. As shown in step of Figure 3, the unlabeled dataset is divided into a WeakFlip set and a StrongFlip set . In step , we perform semi-supervised training using a unified objective that combines supervised contrastive loss with the standard semi-supervised loss, enhancing class separation, particularly for ambiguous and drifted samples. To support adaptation under limited labeling budgets, we design a multi-criteria sample selection strategy in step , which guides the AL process at step , based on the budget defined in step . The cycle from steps to is repeated until the new test set becomes available.
IV-A Augmentations for Malware Representations
CITADEL is built upon FixMatch [22], which is originally developed for the computer vision domain, where augmentations such as rotation, translation, and color distortion are semantically meaningful. In contrast, malware classification operates on a semantically rich, high-dimensional tabular feature space. These features are typically derived from static analysis of binaries (e.g., header metadata, byte sequences, and API call usage), and are subject to strong semantic constraints [47]. Consequently, image-based augmentations are not applicable in this domain and often fail to produce valid perturbations. In the case of Android malware, features are commonly represented as static binary vectors, where each dimension indicates the presence or absence of specific behavioral attributes, permissions, or API calls [6]. These binary vectors encode meaningful structural and behavioral properties of malware samples.
To address this challenge, we propose two augmentation strategies (shown in Algorithm 1 line: 6-7) tailored for static binary feature vectors—Bernoulli bit flip and Bernoulli feature mask (see Figure 4). Since there is no deterministic pattern of distribution shifts caused by concept drift in Android malware [51], we adopt a randomized approach to simulate such drift directly in the feature space through these augmentations. These strategies play a role analogous to data transformations in FixMatch by enabling consistency regularization within the feature space. As a result, the model learns feature representations that remain stable under semantically valid perturbations. MORSE [12] also employs FixMatch, but with a different goal, detecting malware in noisy or incorrectly labeled datasets to improve noise tolerance, without addressing concept drift. Their augmentation method introduces noise by replacing feature values across different samples, where a subset of features is selected according to a Bernoulli distribution.
Bernoulli Bit Flip. This augmentation strategy simulates behavioral mutations commonly observed across malware variants by probabilistically flipping individual bits in the binary feature vector. In this representation, a bit value of 1 indicates the presence of a specific API call or code pattern, while 0 indicates its absence. Given an input feature vector , we initially consider injecting noise using a discrete uniform distribution.
where each is sampled independently and uniformly from , and is the input feature dimension. The noise vector is then combined with the original input using bitwise XOR to generate the augmented sample:
Inspired by Wu et al. [12], we adopt Bernoulli distribution to simulate distribution shift using the original input, as it produces independent binary outcome for each feature with a controlled probability , which aligns with the binary representation of malware features.
Here, denotes the bitwise XOR operation, and is a noise vector sampled independently for each bit, where each bit is flipped with probability . Small values of correspond to subtle modifications that emulate minor obfuscations or mutations, while larger values introduce more substantial perturbations, potentially degrading semantic coherence.
Bernoulli Feature Mask. In this augmentation, we intend to simulate code obfuscation by randomly zeroing out a subset of the features, effectively removing observable behaviors from the sample. Let be a binary mask where each bit is zeroed out with probability . The masked input is defined as:
where denotes element-wise multiplication.
Weak and Strong Augmentation. In CITADEL framework, weak and strong augmentation serve distinct roles. For weak augmentation, we apply a low perturbation probability preserving core semantics of a sample and used for generating confident pseudo-labels, (shown in Algorithm 1 line: 8). For strong augmentation, we apply comparatively high perturbation probability to introduce more noise.
To evaluate the individual contribution of each augmentation strategy, we separately explore the effect of Bernoulli bit flip and Bernoulli feature mask on model generalization and concept drift adaptation. We empirically determine the perturbation rates and to ensure that augmentations are both semantically meaningful and effective for training under concept drift. For Bernoulli bit flip, we apply weak and strong perturbations using flip probabilites and respectively. Similarty, for Bernoulli feature mask, we also apply weak and strong mask with probabilities and , respectively. We also explore a combination of random bit flip and feature masking, and empirically evaluate which perturbation settings provide the best performance under concept drift scenario.
IV-B CITADEL’s Objective Function
We propose an improved objective function incorporating supervised contrastive loss, , computed using the labeled data with labeled samples along with EQ 1 to better distinguish boundary samples and realize effective adaptation to concept drift. This enforces the model to pull together feature representations of samples belonging to the same class while pushing apart those from different classes.
The total loss, (shown in Algorithm 1 line: 15), is defined as a combination of supervised loss and unsupervised consistency loss based on FixMatch, along with the supervised contrastive loss :
where and are weighting hyperparameters for the unsupervised and contrastive terms, respectively. All hyperparameter exploration to select the optimal values of , are provided in Appendix C.
The supervised contrastive loss follows the formulation on [25]. Given a batch of encoded feature vectors and their corresponding labels , the loss is defined as:
where is the temperature scaling parameter that regulates the concentration of the distribution, is the set of indices of positive samples (same label as ) excluding itself. The dot product measures similarity between the features of anchor and positive samples in the representation space. This loss formulation facilitates intra-class compactness (i.e., pulling together samples of the sample class) and inter-class separability (i.e., pushing apart samples of different class) in the representation space. These properties are particularly advantageous in the presence of concept drift where samples near the decision boundary may become ambiguous due to evolving semantic features.
While our semi-supervised framework effectively leverages unlabeled data through pseudo-labeling, it remains vulnerable to real-world drift, where certain samples may be uncertain or misclassified, or poorly represented in the feature space. To further improve adaptability and sample efficiency in such conditions, we integrate an active learning mechanism that selectively queries labels for the most informative and uncertain samples, thereby guiding the model toward better generalization with minimal labeling overhead.
IV-C Multi-Criteria Sample Selection
Active learning offers a promising approach to reducing labeling costs in malware detection [11], particularly in the presence of concept drift. However, selecting informative samples in this context remains challenging due to the inherent sparsity, class imbalance, and evolving nature of malware behaviors. We empirically observe that no single sampling criterion reliably captures all informative instances under such dynamics. To address this, we introduce a multi-criteria sampling strategy based on three complementary indicators: (1) boundary proximity, quantified using the softmax margin; (2) feature displacement, measured via the -norm distance in embedding space; and (3) prediction uncertainty, captured through low-confidence outputs. Each indicator provides a distinct perspective on sample informativeness. We hypothesize that their combination enables more robust and stable adaptation in the face of continual malware evolution. Algorithm 2 presents the detailed procedure of our active learning method.
Softmax Margin for Boundary Samples. The first sampling criterion leverages the model’s softmax prediction margin, which reflects the proximity of a sample to the decision boundary. For an unlabeled input , let denote the model’s softmax output vector. The softmax margin is defined as the difference between the top two predicted class probabilities:
where and are the highest and second-highest values in , respectively. A smaller margin indicates greater uncertainty and typically corresponds to samples near the model’s decision boundary. Such samples are more likely to be informative when labeled. Margin-based uncertainty sampling is a well-established strategy in the active learning literature [52], and has demonstrated robustness even under distributional shift [53].
-Norm for Feature-Space Drift. While the softmax margin captures local uncertainty near the classifier’s decision boundary, it may fail to identify outliers or samples that are semantically distant from the training distribution—a common scenario under concept drift. To address this, we compute the -norm distance between each unlabeled sample and its nearest labeled neighbor in the model’s learned feature space.
Formally, let denote the embedding of a sample obtained from the model’s penultimate layer. For each unlabeled sample and each labeled sample , we compute:
This minimum -distance quantifies the proximity of an unlabeled sample to the closest known (labeled) point in the embedding space. Samples with larger distances are more likely to be unfamiliar or semantically novel, making them strong candidates for labeling under drift.
To select an appropriate value, we empirically evaluated the active learning performance for in increments of 0.1. Our results showed only minor performance variation across values. We ultimately select due to its consistent performance across multiple settings and its theoretical grounding in high-dimensional vector spaces [54, 55].
Low-Confidence Predictions. The third sampling criterion relies on the model’s raw predictive confidence, defined as the maximum probability assigned to any class:
Samples with low confidence often correspond to ambiguous instances or previously unseen malware classes. Prior work has demonstrated that such samples are strongly correlated with misclassification [56] and can be effectively utilized for active learning under distributional shift [36, 35]. In the context of malware detection, this criterion enables the model to prioritize uncertain or suspicious samples that may not be captured by margin-based or distance-based strategies alone.
Sample Selection Strategy. First, we take the intersection of the three criteria to ensure that we do not select the same samples combining the criteria. Then, we leverage the complementary strengths of the three uncertainty criteria: margin-based (), feature-space distance (), and confidence-based (), we compute a unified selection score for each unlabeled sample. As shown in Algorithm 2, lines 3–6, each score component is first min-max normalized across the unlabeled pool:
Here, denotes the normalized value, and are weighting coefficients (set to 1 by default unless otherwise specified). This formulation assigns higher selection scores to samples with low prediction margins, high feature-space distance, and low model confidence—each of which signals potential informativeness under concept drift. The top- samples with the highest scores are selected for labeling.
The sample selection and retraining procedures are detailed in Algorithm 2, lines 10–13, which invoke the semi-supervised training routine in Algorithm 1. Compared to prior active learning approaches in malware detection that rely on a single uncertainty metric [20, 11], our multi-criteria strategy enables more robust adaptation to drift by jointly targeting decision boundary ambiguity and semantic outliers in the learned feature space.
| Year | APIGraph [13] | Chen-AZ [11] | LAMDA [15] | MaMaDroid [14] | ||||
| B | M | B | M | B | M | B | M | |
| 2012 | 27,613 | 3,066 | - | - | - | - | - | - |
| 2013 | 43,873 | 4,871 | - | - | 42,048 | 44,383 | 9,000 | 1,000 |
| 2014 | 52,843 | 5,871 | - | - | 55,427 | 45,756 | 9,000 | 1,000 |
| 2015 | 52,173 | 5,797 | - | - | - | - | - | - |
| 2016 | 50,859 | 5,651 | - | - | 64,059 | 45,134 | 9,000 | 1,000 |
| 2017 | 24,930 | 2,620 | - | - | 77,785 | 21,359 | 9,000 | 1,000 |
| 2018 | 38,214 | 4,213 | - | - | 64,942 | 39,350 | 9,000 | 1,000 |
| 2019 | - | - | 40,947 | 4,542 | 49,465 | 41,585 | 9,000 | 1,000 |
| 2020 | - | - | 34,921 | 3,982 | 55,718 | 46,355 | 9,000 | 1,000 |
| 2021 | - | - | 13,985 | 1,676 | 45,528 | 35,627 | 9,000 | 1,000 |
| 2022 | - | - | - | - | 44,768 | 41,648 | 9,000 | 1,000 |
| 2023 | - | - | - | - | 46,462 | 7,892 | 9,000 | 1,000 |
| 2024 | - | - | - | - | 47,633 | 794 | 9,000 | 674 |
| 2025 | - | - | - | - | 44,640 | 23 | 9,000 | 18 |
| Total | 290,505 | 32,089 | 89,853 | 10,200 | 638,475 | 369,906 | 108,000 | 10,692 |
V Evaluation
In this section, we present the comprehensive evaluation of CITADEL. Section V-A introduces the datasets, followed by Section V-B, which analyzes CITADEL without active learning (AL) under varying labeled ratios and feature augmentations. Section V-C compares CITADEL with baseline methods without AL, while Section V-D extends the comparison to the AL framework. Finally, Section V-E discusses the computational complexity. Additional analysis and experimental results are provided in the Appendix—Appendix A details family and singleton statistics, Appendix B visualizes the performance trends of CITADEL and Chen-AL, Appendix C shows results of hyperparameter tuning, Appendix D examines label noise robustness, Appendix E explores curriculum learning integration, and Appendix F reports results on the EMBER [57].
| LR | Augmentation | APIGraph | Chen-AZ | LAMDA | MaMaDroid | ||||||||
| F1 | FNR | FPR | F1 | FNR | FPR | F1 | FNR | FPR | F1 | FNR | FPR | ||
| 10% | Random Bit Flip | 59.80.7 | 52.90.2 | 7.30.9 | 39.00.6 | 72.01.0 | 0.70.0 | 23.11.3 | 82.50.6 | 3.40.2 | 12.5 0.5 | 92.0 0.8 | 2.8 0.3 |
| Bernoulli Bit Flip | 61.81.4 | 51.41.4 | 5.90.2 | 50.00.5 | 59.70.7 | 1.10.0 | 23.00.4 | 82.50.1 | 3.20.7 | 14.3 0.4 | 90.1 0.6 | 2.6 0.2 | |
| Random Flip + Masking | 60.31.2 | 52.81.4 | 0.60.1 | 37.41.7 | 73.51.2 | 0.60.0 | 22.50.3 | 83.10.0 | 3.00.6 | 13.8 0.4 | 90.7 0.7 | 2.7 0.2 | |
| Bernoulli Masking | 60.31.2 | 52.81.3 | 0.60.1 | 37.70.6 | 74.10.7 | 0.40.1 | 32.11.1 | 68.11.2 | 10.80.3 | 15.8 0.4 | 86.0 0.6 | 3.7 0.3 | |
| 20% | Random Bit Flip | 64.01.6 | 48.82.3 | 6.10.0 | 46.20.3 | 66.01.1 | 0.60.0 | 27.02.1 | 79.03.2 | 3.30.9 | 16.1 0.4 | 87.4 0.6 | 2.6 0.2 |
| Bernoulli Bit Flip | 64.32.0 | 48.22.6 | 0.30.1 | 57.40.0 | 53.50.3 | 0.60.1 | 33.90.2 | 71.31.8 | 4.71.4 | 18.4 0.3 | 84.7 0.5 | 2.4 0.2 | |
| Random Flip + Masking | 64.41.4 | 47.01.1 | 0.90.1 | 46.00.4 | 66.10.5 | 0.50.1 | 29.63.2 | 77.22.6 | 2.50.8 | 17.9 0.3 | 85.2 0.5 | 2.5 0.2 | |
| Bernoulli Masking | 64.41.3 | 47.01.3 | 0.90.1 | 44.90.6 | 66.20.9 | 0.60.1 | 31.70.3 | 75.20.5 | 2.90.2 | 20.4 0.3 | 81.2 0.5 | 3.4 0.2 | |
| 30% | Random Bit Flip | 67.70.6 | 43.60.7 | 8.00.2 | 49.00.7 | 62.70.5 | 0.80.0 | 25.81.1 | 80.51.1 | 2.30.2 | 18.9 0.3 | 82.8 0.5 | 2.5 0.1 |
| Bernoulli Bit Flip | 66.11.8 | 45.43.4 | 7.71.9 | 58.71.0 | 52.11.2 | 0.60.1 | 34.34.0 | 69.54.5 | 4.80.7 | 21.2 0.3 | 79.5 0.4 | 2.3 0.1 | |
| Random Flip + Masking | 66.01.3 | 46.51.5 | 0.60.1 | 47.70.6 | 63.90.7 | 0.70.1 | 27.91.0 | 78.90.8 | 2.40.2 | 20.7 0.3 | 80.0 0.5 | 2.4 0.1 | |
| Bernoulli Masking | 66.01.4 | 46.51.6 | 0.60.1 | 47.71.0 | 63.10.9 | 0.70.1 | 26.73.8 | 78.74.3 | 3.61.3 | 23.6 0.2 | 78.4 0.4 | 3.1 0.2 | |
| 40% | Random Bit Flip | 69.20.9 | 42.52.0 | 6.71.3 | 51.90.5 | 58.90.7 | 0.70.0 | 25.61.0 | 81.01.0 | 1.80.5 | 20.4 0.2 | 79.2 0.4 | 2.4 0.1 |
| Bernoulli Bit Flip | 67.22.9 | 43.63.0 | 9.51.1 | 59.40.7 | 51.30.4 | 0.60.1 | 30.90.9 | 74.80.9 | 2.90.4 | 22.5 0.2 | 76.1 0.3 | 2.2 0.1 | |
| Random Flip + Masking | 70.91.8 | 38.53.1 | 1.00.3 | 51.90.8 | 59.60.9 | 0.50.1 | 24.71.6 | 81.71.6 | 1.90.6 | 22.0 0.2 | 76.6 0.3 | 2.3 0.1 | |
| Bernoulli Masking | 72.21.6 | 36.22.2 | 1.10.2 | 49.81.0 | 61.80.6 | 0.60.1 | 30.73.1 | 75.32.6 | 3.90.5 | 26.2 0.1 | 75.6 0.2 | 2.9 0.0 | |
| 50% | Random Bit Flip | 68.10.4 | 41.10.3 | 11.41.5 | 56.00.5 | 55.70.8 | 0.60.0 | 26.60.0 | 80.10.2 | 2.40.2 | 20.6 0.2 | 78.8 0.4 | 2.4 0.1 |
| Bernoulli Bit Flip | 68.31.9 | 41.82.4 | 8.20.7 | 58.60.7 | 52.60.6 | 0.50.0 | 26.90.6 | 80.00.7 | 2.10.3 | 22.7 0.2 | 75.7 0.3 | 2.2 0.1 | |
| Random Flip + Masking | 68.52.4 | 42.33.2 | 0.70.1 | 52.10.1 | 59.20.0 | 0.60.0 | 28.82.4 | 77.72.6 | 3.10.6 | 22.2 0.2 | 76.2 0.3 | 2.3 0.1 | |
| Bernoulli Masking | 67.21.4 | 44.62.1 | 0.70.1 | 51.60.7 | 59.20.7 | 0.80.2 | 25.70.1 | 80.30.1 | 3.30.3 | 26.8 0.2 | 74.8 0.3 | 2.8 0.1 | |
| 60% | Random Bit Flip | 69.01.9 | 40.80.8 | 10.12.5 | 52.50.4 | 58.80.5 | 0.50.0 | 25.62.4 | 80.63.1 | 2.50.3 | 20.9 0.2 | 78.3 0.4 | 2.3 0.1 |
| Bernoulli Bit Flip | 69.01.6 | 41.02.0 | 9.10.0 | 57.50.2 | 53.40.3 | 0.50.0 | 32.20.5 | 75.60.2 | 1.90.1 | 23.0 0.2 | 75.3 0.3 | 2.1 0.1 | |
| Random Flip + Masking | 66.01.5 | 44.92.2 | 1.00.1 | 50.82.2 | 60.42.2 | 0.50.1 | 21.42.3 | 84.22.0 | 1.80.2 | 22.5 0.2 | 75.8 0.3 | 2.2 0.1 | |
| Bernoulli Masking | 66.01.8 | 44.92.4 | 1.10.5 | 50.60.9 | 60.50.8 | 0.50.1 | 23.11.3 | 82.31.2 | 3.10.6 | 27.2 0.2 | 74.0 0.3 | 2.7 0.1 | |
| 70% | Random Bit Flip | 69.21.0 | 39.70.8 | 10.73.1 | 49.30.6 | 62.11.0 | 0.70.0 | 26.10.6 | 80.80.6 | 2.00.3 | 21.1 0.2 | 77.9 0.3 | 2.3 0.1 |
| Bernoulli Bit Flip | 70.61.1 | 40.91.1 | 7.20.6 | 61.50.0 | 48.80.0 | 0.70.0 | 29.11.2 | 77.70.7 | 2.10.1 | 23.2 0.2 | 75.0 0.3 | 2.1 0.1 | |
| Random Flip + Masking | 67.91.7 | 42.62.5 | 1.00.2 | 56.10.4 | 53.90.9 | 0.70.0 | 22.80.2 | 82.80.8 | 2.50.0 | 22.7 0.2 | 75.5 0.3 | 2.2 0.1 | |
| Bernoulli Masking | 66.91.3 | 44.61.6 | 0.90.3 | 49.00.6 | 61.90.7 | 0.60.0 | 22.40.2 | 82.50.1 | 4.00.3 | 27.6 0.3 | 73.5 0.4 | 2.7 0.1 | |
| 80% | Random Bit Flip | 68.32.2 | 42.22.7 | 8.00.5 | 48.70.6 | 62.70.8 | 0.60.0 | 28.03.2 | 78.73.1 | 2.10.4 | 21.3 0.2 | 77.6 0.3 | 2.3 0.1 |
| Bernoulli Bit Flip | 70.31.6 | 39.62.5 | 8.41.1 | 60.31.0 | 51.01.0 | 0.40.0 | 30.51.0 | 76.30.2 | 2.00.2 | 23.4 0.2 | 74.8 0.3 | 2.1 0.1 | |
| Random Flip + Masking | 69.91.9 | 37.02.7 | 1.70.3 | 49.43.4 | 61.23.5 | 0.70.0 | 24.80.3 | 81.40.3 | 2.20.1 | 22.9 0.2 | 75.3 0.3 | 2.2 0.1 | |
| Bernoulli Masking | 66.11.5 | 45.62.1 | 0.60.1 | 48.70.7 | 63.31.4 | 0.70.1 | 25.01.6 | 80.41.5 | 3.50.1 | 27.9 0.3 | 73.1 0.3 | 2.6 0.1 | |
| 90% | Random Bit Flip | 72.21.5 | 34.54.4 | 13.64.4 | 49.00.6 | 61.61.1 | 0.60.0 | 25.83.3 | 80.33.1 | 2.70.2 | 21.5 0.2 | 77.3 0.3 | 2.3 0.1 |
| Bernoulli Bit Flip | 73.62.4 | 34.94.2 | 10.41.7 | 58.10.7 | 52.11.2 | 0.80.1 | 33.13.4 | 74.23.5 | 2.60.2 | 23.5 0.2 | 74.6 0.3 | 2.1 0.1 | |
| Random Flip + Masking | 70.31.7 | 38.02.1 | 1.30.2 | 50.72.1 | 59.63.4 | 0.80.3 | 22.10.9 | 83.30.7 | 2.50.1 | 23.0 0.2 | 75.0 0.3 | 2.2 0.1 | |
| Bernoulli Masking | 67.01.8 | 41.34.3 | 1.50.9 | 47.21.1 | 62.81.1 | 1.00.1 | 23.50.6 | 79.80.5 | 6.50.3 | 28.2 0.3 | 72.6 0.3 | 2.6 0.1 | |
V-A Datasets
In this work, we utilize four Android malware datasets: (i) APIGraph [40], (ii) Chen-AZ, based on the AndroZoo [58] dataset as used by Chen et al. [11], (iii) LAMDA [15] (iv) MaMaDroid [14].
APIGraph spans 2012–2018 and contains 32,089 malware samples from 1,120 malware families and 290,50 benign samples [13]. Chen-AZ includes 10,200 malware samples across 254 families and 89,853 benign samples from 2019 to 2021[11]. LAMDA is the most recent, largest, and temporally diverse dataset, specifically designed to study longitudinal concept drift. It spans 2013–2025 (excluding 2015) and contains over 1 million APKs, of which 37% are labeled as malware. The benchmark includes 1,008,381 total samples, comprising 369,906 malware and 638,475 benign samples, covering 1,380 unique families [15]. We construct a MaMaDroid [14] dataset that provides a complementary Android malware analysis dataset containing 10,000 APKs per year (9,000 benign and 1,000 malicious), except for 2024 and 2025 where fewer samples are available due to limited data. All APK samples originate from LAMDA but are processed using MaMaDroid’s feature-extraction pipeline, enabling the construction of a MaMaDroid-style dataset that preserves LAMDA’s temporal distribution [14]. Table I summarizes the yearly counts of benign (B), malware (M) across these four malware datasets. Appendix A presents the details the year-wise counts of malware families and singleton samples for each dataset.
| Method | APIGraph | Chen-AZ | LAMDA | MaMaDroid | ||||||||
| F1 | FNR | FPR | F1 | FNR | FPR | F1 | FNR | FPR | F1 | FNR | FPR | |
| TRANSCENDENT | 50.3 0.8 | 61.4 1.6 | 1.1 0.0 | 29.3 0.6 | 78.2 1.6 | 1.3 0.0 | 30.4 1.0 | 75.8 0.7 | 16.4 0.6 | 19.4 0.4 | 83.5 0.3 | 3.9 0.1 |
| CADE | 64.6 1.2 | 39.2 1.2 | 2.5 0.3 | 56.0 1.2 | 50.5 1.5 | 2.1 0.4 | 29.5 0.7 | 76.7 1.6 | 4.0 1.2 | 20.3 0.5 | 82.4 0.4 | 3.6 0.1 |
| Chen-AL | 60.4 1.0 | 47.2 1.3 | 2.0 0.1 | 47.5 0.9 | 61.3 0.5 | 1.6 0.2 | 22.6 0.1 | 83.5 0.1 | 1.7 0.0 | 20.1 0.3 | 81.7 0.4 | 3.2 0.1 |
| FixMatch-Baseline | 0.0 0.0 | 1.0 0.0 | 0.0 0.0 | 0.0 0.0 | 1.0 0.0 | 0.0 0.0 | 0.0 0.0 | 1.0 0.0 | 0.0 0.0 | 0.0 0.0 | 1.0 0.0 | 0.0 0.0 |
| MORSE | 64.5 0.4 | 49.6 0.5 | 0.2 0.0 | 55.5 0.8 | 54.9 0.9 | 0.6 0.0 | 27.1 0.6 | 79.6 0.8 | 1.8 0.0 | 23.2 0.2 | 78.6 0.3 | 2.6 0.1 |
| CITADEL (Ours) | 67.2 2.9 | 43.6 3.0 | 9.5 1.1 | 59.4 0.7 | 51.3 0.4 | 0.6 0.1 | 30.9 0.9 | 74.8 0.9 | 2.9 0.4 | 26.2 0.1 | 75.6 0.2 | 2.9 0.0 |
V-B Evaluating Label Ratios and Feature Augmentations
We evaluate CITADEL under varying supervision levels using multiple label ratios (10%–90%) with our proposed augmentations: random bit flip, random bit flip using Bernoulli, random feature masking, and a combined bit flip and masking strategy.
Experimental Setup. Following the protocol from prior work [11], we adopt the similar multilayer perceptron (MLP) architecture across all experiments to ensure a fair, model-agnostic comparison. For APIGraph, we use 2012 for training, Jan–Jun 2013 for validation, and Jul 2013–Dec 2018 for testing. For Chen-AZ, we use 2019 for training, Jan–Jun 2020 for validation, and Jul 2020–Dec 2021 for testing. LAMDA provides a predefined split–2013 for training, Jan–Jun 2014 for validation, and Jul 2014–Jan 2025 for testing. For MaMaDroid, we adopt the same temporal split as LAMDA for training, validation, and testing. Model performance is reported as the average over all test months. For each label ratio, the training data is randomly divided into labeled and unlabeled subsets, simulating different levels of supervision.
Results. Our results show that the proposed augmentation strategies yield promising performance across varying label ratios. As expected, F1 scores generally improve with increased supervision; however, concept drift still constrains performance, particularly in datasets with severe distributional shifts (see Table II). Among the four datasets, LAMDA exhibits the most significant drift. For instance, with 40% labeled data and the bit flip using Bernoulli augmentation, it achieves an F1 score of around 30% and a false negative rate (FNR) of around 75%. For MaMaDroid dataset with same labeled ratio with Bernoulli masking, it achieves an F1 score of 26% and FNR of 75%. In comparison, APIGraph under the same setting reaches around 70% F1 with a reduced FNR of around 43%. Chen-AZ lies in between, with around 50% F1 and around 50% FNR, reflecting moderate drift. These results highlight that our augmentations, while effective, must be embedded within a continuously updating learning system. Due to the evolving nature of malware, one-time training is insufficient, models must adapt incrementally to temporal concept drift [11].
Across all settings, the Bernoulli bit flip augmentation consistently outperforms others except for MaMaDroid. This suggests that introducing stochastic noise yields more meaningful transformations in binary feature space, enhancing generalization under drift. In addition, we analyzed the impact of the label ratio and observed that performance scores did not improve significantly beyond a 40% ratio. After this point, any performance gains were subtle across all datasets. This finding aligns with established research showing that the benefit of SSL diminishes as the proportion of labeled data grows [22]. Therefore, based on this analysis, we selected the Bernoulli bit flip augmentation with a 40% label ratio for all subsequent experiments.
| Budget | Method | SC | APIGraph | Chen-AZ | LAMDA | MaMaDroid | ||||||||
| F1 | FNR | FPR | F1 | FNR | FPR | F1 | FNR | FPR | F1 | FNR | FPR | |||
| 50 | Binary SVM | UNC | 87.1 0.2 | 17.6 0.4 | 0.7 0.0 | 58.0 0.2 | 53.5 0.4 | 0.2 0.0 | 37.1 0.2 | 5.7 0.3 | 87.9 0.2 | 18.0 0.5 | 77.0 0.7 | 12.4 0.3 |
| Binary GBDT | UNC | 77.8 0.2 | 32.1 0.4 | 0.5 0.0 | 58.9 0.2 | 52.1 0.2 | 0.5 0.0 | 37.0 0.3 | 5.9 0.2 | 87.5 0.3 | 18.3 0.5 | 76.5 0.6 | 12.6 0.3 | |
| TRANSCENDENT | Cred | 75.0 0.3 | 32.0 0.4 | 0.0 0.0 | 38.4 0.3 | 69.4 0.6 | 1.6 0.0 | 32.0 1.5 | 37.8 2.1 | 63.7 1.8 | 19.0 0.5 | 75.5 0.7 | 12.0 0.3 | |
| CADE | OOD | 83.9 1.7 | 19.7 2.0 | 1.0 0.1 | 49.7 0.7 | 62.5 0.4 | 0.3 0.0 | 34.2 1.8 | 44.0 2.5 | 56.8 2.2 | 20.5 0.5 | 73.5 0.6 | 11.4 0.3 | |
| Chen-AL | PS | 89.2 0.3 | 15.1 0.3 | 0.5 0.0 | 69.5 2.2 | 39.6 3.1 | 0.5 0.0 | 37.4 2.4 | 3.5 0.3 | 93.7 2.7 | 22.0 0.4 | 72.0 0.6 | 10.8 0.3 | |
| MORSE | UNC | 88.8 0.3 | 14.9 0.4 | 0.6 0.1 | 55.9 1.2 | 28.2 0.9 | 15.6 0.4 | 61.4 0.4 | 36.1 0.5 | 11.0 0.3 | 25.3 0.9 | 81.2 1.3 | 3.8 0.5 | |
| CITADEL | MC | 90.1 0.1 | 14.5 0.2 | 0.3 0.0 | 72.7 1.0 | 36.7 1.6 | 0.3 0.0 | 70.9 1.2 | 33.6 1.6 | 2.0 0.1 | 30.9 0.9 | 74.8 0.9 | 2.9 0.4 | |
| 100 | Binary SVM | UNC | 87.4 0.3 | 16.6 0.4 | 0.7 0.0 | 59.8 0.2 | 51.5 0.2 | 0.2 0.0 | 37.3 0.3 | 5.4 0.3 | 88.5 0.4 | 20.2 0.4 | 72.7 0.6 | 11.1 0.3 |
| Binary GBDT | UNC | 80.1 0.3 | 27.9 0.3 | 0.7 0.0 | 59.4 0.2 | 50.3 0.3 | 0.7 0.0 | 37.2 0.2 | 5.6 0.4 | 88.2 0.5 | 20.4 0.4 | 72.2 0.6 | 11.0 0.3 | |
| TRANSCENDENT | Cred | 76.8 0.2 | 29.0 0.3 | 0.0 0.0 | 39.8 0.4 | 68.5 0.7 | 1.5 0.0 | 35.8 1.3 | 28.4 1.9 | 70.1 1.6 | 21.3 0.4 | 71.2 0.6 | 10.8 0.3 | |
| CADE | OOD | 87.1 0.5 | 14.4 1.1 | 1.1 0.2 | 50.1 0.1 | 62.3 0.1 | 0.3 0.0 | 37.3 1.6 | 37.2 2.1 | 45.2 2.0 | 24.7 0.4 | 67.5 0.5 | 9.9 0.2 | |
| Chen-AL | PS | 90.7 0.2 | 13.2 0.3 | 0.5 0.0 | 70.7 0.7 | 70.7 0.7 | 0.6 0.0 | 38.5 2.2 | 3.3 0.3 | 93.5 2.6 | 26.3 0.3 | 65.9 0.5 | 9.4 0.2 | |
| MORSE | UNC | 89.4 0.3 | 14.3 0.4 | 0.6 0.1 | 63.2 1.0 | 26.7 0.8 | 10.5 0.3 | 62.3 0.4 | 35.5 0.5 | 11.2 0.3 | 28.4 0.8 | 77.0 1.1 | 3.3 0.4 | |
| CITADEL | MC | 91.9 0.1 | 12.1 0.2 | 0.3 0.0 | 75.7 2.1 | 32.7 2.9 | 0.4 0.1 | 73.8 0.5 | 30.0 0.0 | 3.2 0.4 | 37.6 2.0 | 47.6 1.8 | 18.0 0.9 | |
| 200 | Binary SVM | UNC | 87.9 0.2 | 14.6 0.4 | 0.5 0.0 | 63.5 0.2 | 46.4 0.8 | 0.3 0.0 | 37.7 0.3 | 4.8 0.3 | 89.7 0.7 | 21.6 0.4 | 69.1 0.5 | 9.6 0.2 |
| Binary GBDT | UNC | 82.6 0.2 | 24.7 0.2 | 0.6 0.0 | 63.9 0.2 | 45.0 0.3 | 0.8 0.0 | 37.2 0.2 | 5.6 0.4 | 88.2 0.5 | 21.8 0.4 | 68.7 0.5 | 9.5 0.2 | |
| TRANSCENDENT | Cred | 78.2 0.2 | 27.0 0.3 | 0.0 0.0 | 41.2 0.4 | 67.4 0.6 | 1.3 0.0 | 39.0 1.2 | 21.3 1.5 | 76.3 1.4 | 22.8 0.4 | 67.4 0.5 | 9.2 0.2 | |
| CADE | OOD | 88.9 0.2 | 14.0 0.3 | 0.7 0.0 | 50.2 0.0 | 62.2 0.1 | 0.3 0.0 | 38.5 1.3 | 31.8 1.8 | 38.5 1.6 | 26.3 0.3 | 63.0 0.4 | 8.4 0.2 | |
| Chen-AL | PS | 91.7 0.4 | 11.4 0.6 | 0.4 0.0 | 73.6 1.3 | 33.7 1.2 | 0.5 0.0 | 41.0 1.8 | 2.8 0.3 | 91.2 2.3 | 28.0 0.3 | 61.5 0.4 | 7.8 0.2 | |
| MORSE | UNC | 90.0 0.2 | 13.6 0.3 | 0.5 0.1 | 70.5 0.9 | 25.8 0.7 | 5.3 0.2 | 63.8 0.3 | 34.9 0.4 | 11.5 0.3 | 32.6 1.0 | 72.3 1.0 | 2.9 0.3 | |
| CITADEL | MC | 92.6 0.2 | 10.6 0.2 | 0.3 0.0 | 76.1 0.4 | 31.8 0.6 | 0.3 0.0 | 75.4 0.4 | 26.5 1.1 | 2.8 0.2 | 40.8 1.5 | 37.6 2.1 | 20.4 1.0 | |
| 400 | Binary SVM | UNC | 88.9 0.1 | 12.5 0.2 | 0.8 0.0 | 70.7 0.3 | 38.2 0.8 | 0.4 0.1 | 38.4 0.2 | 3.5 0.3 | 92.0 1.4 | 22.7 0.3 | 65.5 0.4 | 8.1 0.2 |
| Binary GBDT | UNC | 86.4 0.2 | 20.0 0.4 | 0.5 0.0 | 70.0 0.2 | 38.0 0.3 | 0.4 0.0 | 38.3 0.2 | 3.6 0.2 | 91.8 0.9 | 22.9 0.3 | 65.1 0.4 | 8.0 0.2 | |
| TRANSCENDENT | Cred | 78.9 0.2 | 25.1 0.2 | 0.0 0.0 | 43.9 0.4 | 65.4 0.6 | 0.9 0.0 | 40.6 1.1 | 15.6 1.3 | 82.8 1.2 | 24.0 0.3 | 63.8 0.4 | 7.8 0.2 | |
| CADE | OOD | 89.1 0.5 | 13.2 0.5 | 0.8 0.1 | 60.1 0.8 | 51.4 0.7 | 0.4 0.1 | 45.4 1.1 | 59.2 1.5 | 10.1 1.2 | 28.0 0.3 | 59.0 0.4 | 6.9 0.1 | |
| Chen-AL | PS | 92.4 0.2 | 10.2 0.5 | 0.4 0.0 | 75.5 1.6 | 31.4 1.9 | 0.5 0.1 | 43.0 1.6 | 2.5 0.2 | 89.8 2.1 | 30.0 0.2 | 57.0 0.4 | 6.2 0.1 | |
| MORSE | UNC | 90.6 0.2 | 12.9 0.3 | 0.5 0.0 | 78.4 1.1 | 25.1 0.6 | 0.9 0.1 | 64.5 0.3 | 33.8 0.4 | 11.7 0.3 | 36.0 0.7 | 68.1 0.9 | 2.6 0.3 | |
| CITADEL | MC | 93.5 0.3 | 9.5 0.7 | 0.3 0.0 | 82.7 2.6 | 23.9 3.0 | 0.3 0.1 | 77.7 0.1 | 24.0 0.1 | 2.3 0.4 | 44.9 0.0 | 52.7 0.0 | 5.0 0.0 | |
V-C Baseline Comparison without Active Learning
We begin by evaluating CITADEL and prior methods in a static temporal setting, where neither active learning nor model retraining is performed. In this setup, models are trained once on an initial labeled dataset and evaluated on future, temporally shifted data without any incremental updates. This evaluation highlights each model’s ability to generalize under distributional shifts and provides a direct measure of concept drift severity.
Experimental Setup. For the baseline experiments, we follow the same training and evaluation protocol described in Section V-B. In this specific setup, however, no active learning or model updating is applied; models are trained once on the initial labeled data and evaluated on future test sets without any incremental retraining or adaptation.
We compare our system with several representative methods. TRANSCENDENT [10] that uses conformal prediction to reject uncertain predictions without adapting the model. CADE [20] relies on a contrastive autoencoder to detect out-of-distribution (OOD) samples but lacks any adaptation mechanism. Chen-AL [11] applies hierarchical contrastive loss. We also include a FixMatch baseline that uses the original augmentation designed for image classification, which is not optimized for binary malware representations. Furthermore, we evaluate MORSE [12], which applies FixMatch-SSL for malware detection on noisy datasets. n contrast, CITADEL introduces a bit-flip augmentation with a Bernoulli distribution variant, trained at a 40% label ratio. For a fair comparison with MORSE, we adopt the same experimental setup across experiments.
Results. Table III shows that all baseline models suffer substantial performance degradation over time due to concept drift. On the APIGraph dataset, CITADEL achieves 67.2% F1 with only 40% labeled data—outperforming Chen-AL by nearly 7% and reducing the false negative rate (FNR) by over 1.8%. Similar trends appear on Chen-AZ, where CITADEL attains 59.4% F1 which is approximately 12% higher than Chen-AL with a 10% FNR reduction.
LAMDA presents the most challenging setting, reflecting long-term drift from 2014 to 2025. Here, CITADEL achieves 30.9% F1, significantly outperforming Chen-AL by 22.6%, highlighting the limitations of static retraining and supervised contrastive learning. Across all datasets, baselines consistently underperform without active learning, and FixMatch with image-based augmentations performs particularly poorly, reinforcing the importance of malware-specific transformations and adaptive learning strategies under drift scenarios.
These findings emphasize that semi-supervised learning with tailored augmentations is crucial for maintaining performance under drift, especially in low-label regimes. Contrastive learning alone struggles with long-term generalization [11, 20]. To address this, we next introduce an active learning framework that selectively queries the most informative samples under budget constraints, enabling continual adaptation to evolving malware threats.
V-D Comparison with Active Learning
Building on our baseline evaluation, we now incorporate active learning with CITADEL to adapt the semi-supervised model over time in the presence of concept drift. This section presents our active learning setup and compares CITADEL against prior approaches.
Experimental Setup. We follow the same temporal data partitioning protocol described in Section V-B. Consistent with prior work [11], we adopt a warm-start strategy in which the baseline model trained on the initial labeled dataset serves as the starting point. This approach has been shown to provide a more stable foundation for iterative adaptation compared to cold-start methods.
At each test month, we first evaluate the model’s performance on the current data. We then apply our multi-criteria sample selection strategy to identify the most informative unlabeled samples under various labeling budgets (i.e., 50, 100, 200, and 400). The selected samples are annotated and added to the labeled pool, while the remaining test samples are added to the unlabeled pool. The model is subsequently retrained on the updated pools, enabling it to progressively adapt to evolving malware distributions over time.
Active Learning Comparison with Prior Works. We compare our active learning setup with five prior techniques: Binary SVM, Binary GBDT, TRANSCENDENT [29], CADE [20], MORSE [12], and the state-of-the-art Chen-AL [11].
Chen-AL is the first to explicitly design an active learning framework for Android malware detection. Earlier approaches such as CADE [20] is developed for out-of-distribution (OOD) detection using contrastive autoencoders with a rejection mechanism, but they did not incorporate an active learning loop. Similarly, TRANSCENDENT [29] proposes a non-conformity-based criteria to identify and reject harmful samples, enhancing classifier robustness without performing explicit sample selection or iterative labeling. Although MORSE [12] does not address concept drift, its SSL approach for handling noisy malware datasets makes it a relevant baseline for comparison. To ensure fairness, we evaluate MORSE within the same active learning framework by integrating an uncertainty-based sample selection strategy. We include all these methods in our evaluation to benchmark the effectiveness of CITADEL in handling evolving malware distributions.
Results. Table IV presents the comparative performance of CITADEL against prior state-of-the-art approaches across various active learning budgets (50, 100, 200, and 400 samples/month) across the four different datasets. The reported metrics include F1 score, false positive rate (FPR), and false negative rate (FNR) for all four datasets.
Across all datasets and budgets, CITADEL consistently outperforms existing methods in terms of F1, FNR, and FPR. On APIGraph, CITADEL achieves an F1 score of 90.17% with a budget of 50, increasing to 93.52% at a budget of 400. Compared to Chen-AL, our method shows consistent F1 gains (e.g., 1.13% at budgets 100 and 200), while simultaneously reducing FNR and FPR, demonstrating improved handling of uncertain and evolving malware samples. In contrast, CADE and TRANSCENDENT, which rely on OOD rejection without model adaptation, yield lower F1 scores and higher FNRs, particularly under lower labeling budgets.
The performance gap is even more pronounced on the Chen-AZ dataset. At a budget of 400, CITADEL achieves an F1 score of 82.69%, outperforming Chen-AL by 7.12% and reducing the FNR from 31.43% to 23.95%. Even with only 50 labeled samples per month, CITADEL maintains a strong F1 of 72.76%, demonstrating greater label efficiency than competing methods. These gains highlight the system’s ability to generalize under moderate concept drift and noisy unlabeled data. LAMDA and the curated MaMaDroid datasets, spanning over a decade, presents the most challenging evaluation due to their severe and long-term distributional drift. In these datasets, CITADEL achieves F1 scores of 77.6% 44.9% on LAMDA and MaMaDroid, respectively, at a budget of 400—an improvement of over 34% and 14.9% compared to Chen-AL (43.0% on LAMDA and 30% on MaMaDroid). While Chen-AL reports a lower FNR of 2.5%, it suffers from an extremely high FPR of 90%, indicating poor generalization. In contrast, CITADEL delivers balanced performance, with an FNR of 24.05% and a remarkably low FPR of 2.35%, demonstrating its robustness and adaptability in realistic, evolving malware detection scenarios.
On the LAMDA dataset, MORSE achieves an F1 score of 64.5%, with an FNR of 33.8%. On MaMaDroid, its performance drops further to an F1 of 36.0%, FNR of 68.1%. In contrast, CITADEL attains substantially higher F1 scores on LAMDA and on MaMaDroid, while also maintaining lower FNR and FPR values. This comparison highlights that CITADEL demonstrates superior performance over MORSE, primarily due to its more effective augmentations, objective function and sample selection techniques.
| n | Chen-AL (S) | CITADEL (S) | Speedup | Chen-AL Op | CITADEL Op | OpRed |
| 210.0 | 1.7 | 123.5 | 2.63e6 | 3.89e4 | 67.6 | |
| 118.0 | 1.0 | 118.0 | 3.14e6 | 4.31e4 | 72.9 | |
| 188.0 | 6.1 | 30.8 | 5.35e6 | 2.13e5 | 25.1 | |
| 273.0 | 10.2 | 26.8 | 8.11e6 | 4.25e5 | 19.1 | |
| 1030.0 | 45.4 | 22.7 | 3.02e7 | 2.12e6 | 14.2 | |
| 2312.0 | 95.8 | 24.1 | 5.79e7 | 4.24e6 | 13.6 | |
| 9909.0 | 464.4 | 21.3 | 2.58e8 | 2.12e7 | 12.2 |
| Configuration | APIGraph | MaMaDroid | LAMDA | ||||||
| F1 Score | FNR | FPR | F1 Score | FNR | FPR | F1 Score | FNR | FPR | |
| Baseline and Augmentation Only | |||||||||
| FixMatch-Baseline (No Aug, No AL) | 0.00.0 | 1.00.0 | 0.00.0 | 0.00.0 | 1.00.0 | 0.00.0 | 0.00.0 | 1.00.0 | 0.00.0 |
| + Our augmentation only (no AL) | 67.22.9 | 43.63.0 | 9.51.1 | 26.2 0.1 | 75.6 0.2 | 2.9 0.0 | 30.90.9 | 74.80.9 | 2.90.4 |
| Baseline + Augmentation + Active Learning | |||||||||
| + AL (Random Sampling) | 79.91.5 | 25.42.0 | 1.20.3 | 27.8 0.1 | 76.6 0.2 | 2.9 0.0 | 36.13.8 | 67.24.2 | 5.20.6 |
| + AL (Margin Only) | 92.30.2 | 11.00.3 | 0.40.0 | 42.7 0.1 | 54.1 0.2 | 5.3 0.0 | 75.71.4 | 25.71.2 | 2.00.2 |
| + AL (-Norm Only) | 92.10.2 | 12.10.2 | 0.30.0 | 43.1 0.9 | 53.7 0.7 | 5.2 0.3 | 75.30.2 | 26.80.3 | 2.40.2 |
| + AL (Low Confidence Only 75%) | 85.10.2 | 20.10.3 | 0.50.0 | 28.4 0.1 | 73.9 0.2 | 2.6 0.0 | 69.10.2 | 35.90.3 | 2.10.2 |
| Baseline + Augmentation + Active Learning With Multi-Criteria Selection | |||||||||
| + Our Multi-Criteria (w/o Loss) | 93.00.2 | 10.20.2 | 0.30.0 | 44.6 0.6 | 53.0 0.5 | 5.2 0.3 | 77.00.2 | 24.80.2 | 2.60.3 |
| + Our Multi-Criteria + Combined Loss | 93.50.3 | 9.50.7 | 0.30.0 | 44.9 0.9 | 52.7 0.8 | 5.0 0.4 | 77.70.1 | 24.00.1 | 2.30.4 |
To ensure a fair and credible comparison, we re-evaluated Chen-AL under strict temporal partitioning, with all validation samples excluded from the active learning retraining process. These experiments are conducted independently by multiple researchers, each repeated over five runs. Final results are reported as the mean ± standard deviation. Our re-evaluation yields average F1 scores of 75.57% on the Chen-AZ dataset and 92.39% on the APIGraph dataset.
V-E Complexity Analysis
We evaluate the computational scalability of CITADEL in comparison to prior methods. Using a fixed batch size of 10, input dimensionality, and training budget of 400 samples, we measure runtime and operation count over one epoch of initial training and one epoch of active learning retraining. Experiments are conducted on random subsamples from the APIGraph dataset with increasing sizes ().
To quantify computational cost, we use PyTorch Profiler to record total runtime and number of operations across all methods. As shown in Figure 5(a), CITADEL consistently maintains significantly lower runtime across increasing dataset sizes. Figure 5(b) further illustrates that CITADEL requires substantially fewer operations. At , CITADEL completes both training and retraining epochs in just 95.8 seconds and requires approximately 4.2 million operations. In comparison, Chen-AL [11] takes 2,312 seconds and performs 57.9 million operations for the same task. This results in a 24.1 reduction in runtime and a 13.7 reduction in operation count, demonstrating the computational efficiency of CITADEL. CADE [20] incurs comparable overhead to Chen-AL due to clustering and latent-space projections, while TRANSCENDENT [29] is even slower, particularly at scale and in high-dimensional input spaces.
All detailed comparisons of runtime and operation counts between CITADEL and Chen-AL are summarized in Table V. This level of efficiency allows our system to adapt more quickly to emerging threats and deliver more timely updates to end users in realistic deployment environments.
VI Ablation Study
We conduct an ablation study to assess the contribution of individual components within CITADEL. First, we evaluate the FixMatch [22] baseline without our feature augmentations. Then, we incorporate our proposed augmentation strategy (Bernoulli bit flip) and evaluate the model without any active learning using 40% labeled ratio. To analyze the effectiveness of our multi-criteria sample selection method, we compare it against several alternative active learning strategies for 400 budgets, including selection based solely on decision boundary sample selector [53, 52] (softmax margin), -norm uncertainty sample selector, low-confidence samples, and random sample selector. Finally, we evaluate the contribution of CITADEL objective function, which explicitly helps the separation of boundary samples in the latent space. The results demonstrate that each component of CITADEL contributes to improved adaptation under concept drift.
We conducted the ablation study on three datasets–APIGraph [13], MaMaDroid [14] and LAMDA [15]. LAMDA and Chen-AZ are both DREBIN [6] feature-based and collected from AndroZoo repository [58]. In contrast, APIGraph contains structured API call graph features and MaMaDroid utilizes abstraction. Therefore, we choose these three dataset because of their difference in feature representation.
Table VI show each component contributes to performance improvements on the three datasets APIGraph, MaMaDroid and LAMDA. By applying our tailored augmentation techniques, the model achieves improved performance of 67.2% on APIGraph and 30.9% on LAMDA. Building on this, the addition of our muti-criteria sample selection strategy with CITADEL objective function further enhances the model’s ability to adapt to concept drift. Overall, CITADEL achieves final F1-scores of 93.5%, 44.9%, and 77.7% on APIGraph, MaMaDroid, and LAMDA respectively.
VII Discussion
While CITADEL achieves strong performance under concept drift, it remains vulnerable to evasion and poisoning attacks like all ML-based malware detectors. Reliance on pseudo-labels and selective labeling introduces risks from clean-label poisoning, where crafted samples may gradually influence decision boundaries. Although the multi-criteria sample selection helps reduce these effects, additional research is required to strengthen its robustness against adversarial labeling and evasion in future.
Since our method employs continuous drift adaptation, it may introduce a new attack surface. For instance, an adversary could monitor the adaptive behavior of the malware detector using black-box API queries to train a surrogate (proxy) detector. Once the proxy is trained, the attacker can use it to craft evasive malware samples that are likely to evade detection by aligning with the decision boundaries of the target model. We plan to address this type of adversarial vulnerability in future work.
Our findings demonstrate that semi-supervised learning with active sample selection is a viable alternative to fully supervised methods for malware detection in drift-heavy environments. Unlike existing SOTA [11, 20] methods that rely on semantic structures (e.g., API graphs) within malware, our approach is designed to remain effective even in the absence of such structures, particularly under long-term evaluation scenarios. Further, CITADEL consistently performs well even with limited labels and achieves high computational efficiency. However, performance gains taper off beyond a 40% label ratio, and classification remains challenging for singleton and ambiguous samples. Contrastive loss improves boundary separation, but additional techniques like denoising or adaptive representation learning may be beneficial.
VIII Conclusion
In this work, we introduce CITADEL, a semi-supervised active learning framework for Android malware detection under long-term concept drift. CITADEL combines malware domain focused augmentation strategies tailored to binary feature vectors with a multi-criteria active sample selection mechanism. Through extensive evaluations on four real-world Android malware datasets, we demonstrate that CITADEL consistently outperforms state-of-the-art baselines. Notably, on the most temporally drifted LAMDA dataset, CITADEL achieves up to a 14% improvement in F1 score while requiring only 40% of the labeled data. It also delivers substantial efficiency gains, with up to faster training and fewer operations. We believe that semi-supervised and active learning-based approaches like CITADEL hold strong promise for real-world deployment due to their robustness, efficiency, and adaptability in the face of evolving malware threats.
Appendix A Family and Singleton Statistics in the Datasets
| Dataset | 2012 | 2013 | 2014 | 2015 | 2016 | 2017 | 2018 | 2019 | 2020 | 2021 | 2022 | 2023 | 2024 | 2025 | ||||||||||||||
| F | S | F | S | F | S | F | S | F | S | F | S | F | S | F | S | F | S | F | S | F | S | F | S | F | S | F | S | |
| APIGraph [13] | 104 | 36 | 172 | 68 | 175 | 55 | 193 | 53 | 199 | 68 | 147 | 48 | 128 | 45 | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| Chen-AZ [11] | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 121 | 42 | 82 | 31 | 51 | 17 | - | - | - | - | - | - | - | - |
| LAMDA [15] | - | - | 213 | 1320 | 231 | 2046 | - | - | 375 | 4819 | 207 | 7324 | 373 | 16531 | 635 | 15396 | 588 | 24791 | 295 | 24083 | 651 | 20147 | 224 | 4839 | 64 | 538 | 8 | 18 |
| MaMaDroid [14] | - | - | 1 | 32 | 1 | 47 | - | - | 96 | 132 | 48 | 391 | 80 | 521 | 134 | 291 | 117 | 455 | 62 | 805 | 130 | 515 | 93 | 717 | 57 | 526 | 7 | 11 |
Table VII presents the year-wise counts of singleton malware samples—those appearing only once—for the same datasets. APIGraph reports singletons from 2012–2018, peaking at 68 in 2013, with no data afterward. Chen-AZ includes singleton counts from 2019–2021, peaking at 42 in 2019. LAMDA shows extensive singleton data from 2013–2025, with a major surge between 2018 and 2022, peaking at 24,791 in 2020. Dashes indicate years without data. These singleton trends reflect temporal variations and dataset-specific dynamics, potentially signaling malware evolution or labeling noise. Additionally, Table VII also summarizes the year-wise counts of malware families across the four datasets. APIGraph lists family totals from 2012–2018, increasing from 104 families in 2012 to a peak of 199 in 2016, followed by a gradual decline. Chen-AZ provides family counts only for 2019–2021, ranging from 51 to 121 during this period. LAMDA’s family counts span from 2013 to 2025, starting with 213 families in 2013 and increasing over the years, reaching as high as 651 families in 2022. MaMaDroid reports family data for the same years as LAMDA, but its family count is relatively less. It reached a maximum of 134 families in a single year, and its lowest is one family in 2013–2014. Among all datasets, LAMDA exhibits the broadest temporal coverage and the highest family diversity.
Appendix B CITADEL vs Chen-AL performance visualization
Figures 6, 7, 8, and 9 visualize monthly trends in F1-score, false positive rate (FPR), and false negative rate (FNR) for CITADEL versus Chen-AL [11] on the APIGraph, Chen-AZ (AndroZoo), LAMDA, and MaMaDroid datasets, respectively.
Appendix C CITADEL Hyperparameter tuning
We conducted the hyperparameter tuning experiments. First, we conducted the different label ratio experiment to show the reason for selecting 40% labeled ratio for active learning scenario. We saw the performance plateau after 40% labeled ratio.
| Labeled Ratio (%) | Mean F1 | Mean FNR | Mean FPR |
| 10 | 92.990.15 | 9.680.12 | 0.390.05 |
| 20 | 93.090.12 | 9.490.10 | 0.410.04 |
| 30 | 93.390.14 | 9.150.09 | 0.380.05 |
| 40 | 93.570.10 | 8.640.08 | 0.400.04 |
| 50 | 93.600.12 | 8.790.09 | 0.370.05 |
| 60 | 93.580.11 | 8.810.08 | 0.360.04 |
| 70 | 93.450.13 | 9.160.09 | 0.360.05 |
| 80 | 93.470.10 | 9.080.09 | 0.350.05 |
| 90 | 93.510.12 | 8.950.10 | 0.350.04 |
To analyze the impact of augmentation probability on performance, we conducted a systematic tuning experiment over both weak and strong augmentations in CITADEL. Specially, we employ random bit flip from Bernoulli distribution with probability to determine the impact of feature perturbation and fine-tune to get optimal values. We evaluate three settings: (i) fixing the strong augmentation probability () while varying the weak augmentation ; (ii) fixing the weak augmentation probability () while varying the strong augmentation ; and (iii) applying the same augmentation probability for both weak and strong augmentation. As shown in Table IX, the performance remains stable when is small (e.g., 1–5%), but gradually degrades as increases beyond this range.
| Applied aug () | 0.01 | 0.05 | 0.1 | 0.2 | 0.3 | 0.4 | 0.5 | 0.6 | 0.7 |
| Strong Fixed () | 93.5 | 93.5 | 93.4 | 93.3 | 93.1 | 93.1 | 92.8 | 92.1 | 91.7 |
| Vary Weak () | 92.8 | 93.5 | 93.4 | 93.3 | 93.2 | 93.3 | 93.1 | 92.6 | 92.5 |
| Same Weak and Strong () | 93.2 | 93.5 | 93.4 | 93.3 | 93.2 | 93.3 | 93.1 | 92.4 | 91.8 |
While CITADEL adopts several hyperparameters from FixMatch (e.g., and ) as initial references, we further fine-tuned these and introduced a new hyperparameter, . Our experiments show that both and are sensitive to model stability—decreasing or increasing introduces additional noise from unreliable pseudo-labels, thereby degrading performance when learning from unlabeled data. The selected configuration (, ) achieves the best trade-off between leveraging unlabeled data and maintaining detector reliability, confirming that our final settings are empirically justified for this domain. In addition, CITADEL introduces two additional hyperparameters, and , whose optimal values were determined through extensive experimental validation for the malware specific-domain.
Appendix D CITADEL performance under label noise
Experiments. To evaluate the robustness of CITADEL against noisy supervision, we conduct a controlled experiment using the APIGraph dataset. We start with cleanly labeled Android malware samples and progressively inject synthetic label noise by randomly flipping a portion of the ground-truth labels to incorrect ones. The label noise rate varies from 0% to 90% in increments of 10%. The training data may contain misclassified or mislabeled samples, which is common in large-scale malware datasets due to incorrect labels provided by automated anti-virus engines.
Results. CITADEL demonstrates a strong tolerance to label corruption. As shown in Table X, CITADEL maintains an F1-score above 88% for up to 20% label noise, and even with extreme corruption levels of 90% noisy labels, it achieves an F1-score of approximately 70%. This corresponds to only a 23% relative drop compared to the noise-free case. These results highlight that CITADEL retains stable, even under heavily corrupted supervision.
| Noise (%) | F1 | FNR | FPR |
| 0 | 93.5 0.3 | 8.4 0.3 | 0.4 0.1 |
| 10 | 91.3 0.3 | 12.4 0.4 | 0.4 0.1 |
| 20 | 87.8 0.4 | 14.0 0.5 | 0.9 0.1 |
| 30 | 84.8 0.4 | 15.5 0.5 | 1.5 0.2 |
| 40 | 81.5 0.5 | 17.0 0.6 | 2.5 0.3 |
| 50 | 78.1 0.6 | 18.4 0.6 | 4.3 0.4 |
| 60 | 76.2 0.6 | 19.5 0.6 | 6.9 0.5 |
| 70 | 74.3 0.7 | 21.5 0.7 | 9.8 0.6 |
| 80 | 73.6 0.7 | 14.4 0.8 | 11.7 0.7 |
| 90 | 70.0 0.8 | 16.6 0.8 | 13.1 0.8 |
Appendix E CITADEL with curriculum learning
Experiments. To evaluate whether Curriculum Learning (CL) can mitigate malware drift, we designed a baseline experiment where training samples were ordered from easy to hard following the CL principle proposed by [44]. Unlike Active Learning (AL), which adaptively selects new informative samples during training, CL assumes a smooth progression in task difficulty. This assumption rarely holds in malware detection, where new variants appear abruptly. We used the 2012 training subset of the APIGraph dataset and ranked all instances by instance hardness following [44]. To simulate limited human annotation effort, we additionally labeled 400 hard samples per month as in [59]. Both settings used identical hyperparameters and data splits as our main experiments for fair comparison with active learning.
| Method | F1 | FNR | FPR |
| w/o CurL | 71.20.6 | 36.20.7 | 1.200.0 |
| CurL + 400 samples | 83.00.5 | 24.00.6 | 0.900.0 |
| AL (CITADEL) | 93.50.3 | 8.40.3 | 0.40.1 |
Results. Table XI summarizes the comparison. Without CL, the model achieved an F1 score of 71.2%, with relatively high FNR and FPR. Incorporating CL and limited hard-sample feedback improved the F1 score to 83.0%, showing moderate gains in generalization and convergence stability. However, our proposed Active Learning (CITADEL) approach—jointly considering margin, distance, and confidence (Table IV)—achieved a substantially higher 93.5% F1, with a much lower FNR of 8.4% and FPR of 0.4%. These results demonstrate that while CL helps on static datasets, AL is far more effective under evolving malware conditions due to its adaptability and targeted sample selection.
Appendix F Experiments on EMBER Feature
Experiments. On the EMBER dataset, we use samples from 2018-01 to 2018-02 for initial training, and begin active learning from 2018-03 to 2018-12 using different methods—Chen-AL [11], CADE [20], and CITADEL. For this feature space, since Bernoulli bit-flip is not applicable, we employ Bernoulli feature masking as the augmentation strategy in CITADEL. This approach regularizes the model by randomly masking input features, thereby improving generalization under limited labeling budgets.
Results. Table XII presents the performance of CADE, Chen-AL, and CITADEL under different labeling budgets on the EMBER dataset. CITADEL consistently achieves the best results across all budgets. At a budget of 50, CITADEL attains an F1-score of 87.9%, outperforming CADE (85.4%) and Chen-AL (83.4%). With a budget of 400, CITADEL further improves to an F1-score of 89.4%, while CADE reaches 86.0%.
| Budget | Method | F1 | FNR | FPR |
| 50 | CADE | 85.40.5 | 22.50.4 | 4.00.2 |
| Chen-AL | 83.40.8 | 17.30.5 | 18.50.6 | |
| CITADEL | 87.90.5 | 16.80.4 | 6.60.3 | |
| 100 | CADE | 85.70.5 | 22.10.4 | 4.00.2 |
| Chen-AL | 83.70.8 | 17.00.5 | 18.10.5 | |
| CITADEL | 88.40.5 | 16.10.4 | 6.40.3 | |
| 200 | CADE | 85.90.4 | 21.80.3 | 4.10.2 |
| Chen-AL | 84.00.7 | 16.70.4 | 17.60.5 | |
| CITADEL | 88.90.4 | 15.30.3 | 6.30.2 | |
| 400 | CADE | 86.00.4 | 21.50.3 | 4.20.2 |
| Chen-AL | 84.30.4 | 16.50.3 | 16.90.3 | |
| CITADEL | 89.40.3 | 14.50.3 | 6.30.2 |